# Rectifying Magnitude Neglect in Linear Attention

### v1

@@ Abstract
As the core operator of Transformers, Softmax Attention exhibits excellent global modeling capabilities. However, its quadratic complexity limits its applicability to vision tasks. In contrast, Linear Attention shares a similar formulation with Softmax Attention while achieving linear complexity, enabling efficient global information modeling. Nevertheless, Linear Attention suffers from a significant performance degradation compared to standard Softmax Attention. In this paper, we analyze the underlying causes of this issue based on the formulation of Linear Attention. We find that, unlike Softmax Attention, Linear Attention entirely disregards the magnitude information of the Query($Q$ or $\phi(Q)$). This prevents the attention score distribution from dynamically adapting as the Query scales. As a result, despite its structural similarity to Softmax Attention, Linear Attention exhibits a significantly different attention score distribution. Based on this observation, we propose Magnitude-Aware Linear Attention (MALA), which modifies the computation of Linear Attention to fully incorporate the Query’s magnitude. This adjustment allows MALA to generate an attention score distribution that closely resembles Softmax Attention while exhibiting a more well-balanced structure. We evaluate the effectiveness of MALA on multiple tasks, including image classification, object detection, instance segmentation, semantic segmentation, natural language processing, speech recognition, and image generation. Our MALA achieves strong results on all of these tasks. Code will be available at https://github.com/qhfan/MALA.
@@ 4 Experiments
Cost Model Type Parmas (M) FLOPs (G) Top1-acc (%)
Tiny model $\sim 2.5$G NAT-M [23] Trans 20 2.7 81.8
@@
Small model $\sim 4.5$G MogaNet-S [32] CNN 25 5.0 83.4
@@
Base model $\sim 10.0$G ConvNeXT-S [38] CNN 50 8.7 83.1
@@
Large model $\sim 15.0$G MogaNet-L [32] CNN 83 15.9 84.7
@@
Table 2: Comparison with the state-of-the-art on ImageNet-1K classification. We use "CNN" to refer to convolutional neural networks, "Trans" to refer to Vision Transformers, "Mamba" to refer to visual state space model, and "Linear" to refer to models based on Linear Attention.
Backbone Type Params (M) FLOPs (G) $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
@@
Table 3: Comparison to other backbones on "$3\times+\mathrm{MS}$" schedule.
Backbone Type Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
@@
Table 4: Comparison to other backbones with “1×1$\times$1 ×" schedule.
@@ 4.2 Object Detection and Instance Segmentation
Settings. Following previous works [13, 37, 57], we use RetinaNet [33], Mask-RCNN [25] and Cascade Mask R- CNN [5] to evaluate our models. we use the commonly used "1×1$\times$1 ×" (12 epochs) setting for the RetinaNet and Mask R-CNN and "3×3$\times$3 ×+MS" (36 epochs) for Mask R-CNN and Cascade Mask R-CNN.
@@ 4.3 Semantic Segmentation
Settings. Follow previous works [37, 46, 18], we adopt SemanticFPN [29] and UperNet [54] to evaluate our models. For SemanticFPN, we train the models for 80K iterations [51, 52], while for UperNet, we train them for 160K iterations. The batch sizes are set to 16 for all models. The input images are cropped to $512\times 512$.
@@
Params (M) FLOPs (G) mIoU (%) Params (M) FLOPs (G) mIoUss (%)
@@ 4.4 Inference Efficiency
Settings. To thoroughly assess the inference efficiency of MAViT, we evaluate its performance on both low-resolution task (image classification with a resolution of $224\times 224$) and high-resolution task (semantic segmentation with a resolution of $512\times 2048$). For the low-resolution task, we measure model throughput using a batch size of 64. For the high-resolution task, we evaluate the frame per second (FPS) performance with a batch size of 1.
Figure 5: Comparison of general backbones’ inference speed on low resolution task (image classification, resolution $224\times 224$). The inference speed are measured on A100, batch size 64.
Figure 6: Comparison of general backbones’ inference speed on high resolution task (semantic segmentation with UperNet, resolution $512\times 2048$). The inference speed are measured on A100, batch size 1.

### v3

@@ Abstract
As the core operator of Transformers, Softmax Attention exhibits excellent global modeling capabilities. However, its quadratic complexity limits its applicability to vision tasks. In contrast, Linear Attention shares a similar formulation with Softmax Attention while achieving linear complexity, enabling efficient global information modeling. Nevertheless, Linear Attention suffers from a significant performance degradation compared to standard Softmax Attention. In this paper, we analyze the underlying causes of this issue based on the formulation of Linear Attention. We find that, unlike Softmax Attention, Linear Attention entirely disregards the magnitude information of the $Q$ or $\phi(Q)$). This prevents the attention score distribution from dynamically adapting as the Query scales. As a result, despite its structural similarity to Softmax Attention, Linear Attention exhibits a significantly different attention score distribution. Based on this observation, we propose Magnitude-Aware Linear Attention (MALA), which modifies the computation of Linear Attention to fully incorporate the Query’s magnitude. This adjustment allows MALA to generate an attention score distribution that closely resembles Softmax Attention while exhibiting a more well-balanced structure. We evaluate the effectiveness of MALA on multiple tasks, including image classification, object detection, instance segmentation, semantic segmentation, natural language processing, speech recognition, and image generation. Our MALA achieves strong results on all of these tasks. Code will be available at https://github.com/qhfan/MALA.
@@ 4 Experiments
Cost Model Type
Parmas
FLOPs
Top1-acc
Tiny model 2.5$\sim$ 2.5∼ 2.5G
NAT-M [23] Trans 20 2.7 81.8
@@
Small model 4.5$\sim$ 4.5∼ 4.5G
MogaNet-S [32] CNN 25 5.0 83.4
@@
Base model 10.0$\sim$ 10.0∼ 10.0G
ConvNeXT-S [38] CNN 50 8.7 83.1
@@
Large model 15.0$\sim$ 15.0∼ 15.0G
MogaNet-L [32] CNN 83 15.9 84.7
@@
Table 2: Comparison with the state-of-the-art on ImageNet-1K classification. We use ”CNN” to refer to convolutional neural networks, ”Trans” to refer to Vision Transformers, ”Mamba” to refer to visual state space model, and ”Linear” to refer to models based on Linear Attention.
Backbone Type
Params
FLOPs
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
@@
Table 3: Comparison to other backbones on ”$3\times+\mathrm{MS}$” schedule.
Backbone Type
Params
FLOPs
Mask R-CNN 1×1$\times$1 ×
Params
FLOPs
RetinaNet 1×1$\times$1 ×
@@
Table 4: Comparison to other backbones with “1×1$\times$1 ×” schedule.
@@ 4.2 Object Detection and Instance Segmentation
Settings. Following previous works [13, 37, 57], we use RetinaNet [33], Mask-RCNN [25] and Cascade Mask R- CNN [5] to evaluate our models. we use the commonly used ”1×1$\times$1 ×” (12 epochs) setting for the RetinaNet and Mask R-CNN and ”3×3$\times$3 ×+MS” (36 epochs) for Mask R-CNN and Cascade Mask R-CNN.
@@ 4.3 Semantic Segmentation
Settings. Follow previous works [37, 46, 18], we adopt SemanticFPN [29] and UperNet [54] to evaluate our models. For SemanticFPN, we train the models for 80K iterations [51, 52], while for UperNet, we train them for 160K iterations. The batch sizes are set to 16 for all models. The input images are cropped to $\times 512$ 512.
@@
Params
FLOPs
mIoU
Params
FLOPs
mIoUss
@@ 4.4 Inference Efficiency
Settings. To thoroughly assess the inference efficiency of MAViT, we evaluate its performance on both low-resolution task (image classification with a resolution of $\times 224$ 224) and high-resolution task (semantic segmentation with a resolution of 512×2048512$\times$ 2048512 × 2048). For the low-resolution task, we measure model throughput using a batch size of 64. For the high-resolution task, we evaluate the frame per second (FPS) performance with a batch size of 1.
Figure 5: Comparison of general backbones’ inference speed on low resolution task (image classification, resolution $\times 224$ 224). The inference speed are measured on A100, batch size 64.
Figure 6: Comparison of general backbones’ inference speed on high resolution task (semantic segmentation with UperNet, resolution 512×2048512$\times$ 2048512 × 2048). The inference speed are measured on A100, batch size 1.